# Context-aware Biases for Length Extrapolation

### v2

@@ Abstract
Transformers often struggle to generalize to longer sequences than those seen during training—a limitation known as length extrapolation. Most existing Relative Positional Encoding (RPE) methods attempt to address this by introducing either fixed linear biases or globally learned biases, which lack the capacity to adapt to different input contexts. In this work, we propose an additive RPE, Context-Aware Biases for Length Extrapolation (CABLE), a method that learns token-specific, context-aware biases for each attention head in transformers. By dynamically adjusting positional biases based on the input sequence, CABLE overcomes the rigidity of fixed RPEs. When evaluated on sequences longer than originally trained with, GPT-2 Medium (334M parameters) with CABLE achieves lower perplexity than counterparts using other widely adopted positional encoding methods. Additionally, by applying CABLE to the BERT base model we improved performance in long-context retrieval tasks. Our method significantly enhances the extrapolation performance of existing RPE methods tested on the FineWeb-Edu-10B and WikiText-103 datasets. Code is available at: https://github.com/axiomlab/cable.
@@ 2 Related Work
The attention mechanism in Transformer models is inherently position-agnostic, necessitating the explicit injection of positional information to enable sequence order awareness. To address this, a range of positional encoding methods have been proposed. However, many of these methods face limitations when it comes to length extrapolation—the ability to generalize to sequences longer than those encountered during training Anil et al. (2022). Several positional encoding strategies have been developed with the goal of improving extrapolation performance Csordás et al. (2021), with relative positional encodings (RPEs) demonstrating notable effectiveness in enhancing generalization at inference time Song and Zhong (2023); Qin et al. (2024); Dong et al. (2024). In this section, we review key approaches to positional encoding, including absolute and relative methods, as well as recent work exploring Transformer models without any explicit positional encoding.
@@
Fire Li et al. (2023). The kernel function is defined as , where is a monotonically increasing function and is a learnable scaler. This formulation allows Fire to assign more attention to distant query-key pairs—contrary to methods like ALiBi, and Kerple, which tend to focus on nearby tokens. However, as our experiments demonstrate, this behavior was not beneficial in our settings, leading to inferior performance compared to ALiBi and Kerple.
@@
Examining existing additive RPEs, we observe that their bias functions are static—dependent only on token positions and entirely independent of the input context. Once trained, these biases remain fixed across all sequences, limiting their adaptability. This rigidity can hurt performance, especially in long-context generation. To address this, we introduce CABLE, a context-aware additive RPE designed for better length extrapolation in transformer-based language models. Unlike methods such as ALiBi that use fixed linear biases, CABLE learns dynamic, token-specific biases conditioned on the input, enabling more flexible and effective positional encoding.
@@ 4.3 Baselines
RoPE Su et al. (2024): A non-learnable RPE widely adopted in LLMs such as GPT-2 Brown et al. (2020), LLaMA Touvron et al. (2023), PaLM Chowdhery et al. (2023), and Gemma Team et al. (2024a, b).
ALiBi Press et al. (2021): A non-learnable RPE used in models like BLOOM Le Scao et al. (2023) and Falcon Almazrouei et al. (2023).
T5-bias Raffel et al. (2020): A learnable RPE used in the T5 model.
Kerple Chi et al. (2022a): A learnable RPE with logarithmic and power variants; we use the logarithmic variant due to its superior performance.
Fire Li et al. (2023): A learnable RPE designed to give more weight to distant query-key pairs than other methods.
@@ 5 Results
We evaluate the performance of our method across multiple scenarios. First, we assess its extrapolation ability in decoder-only models for next-token prediction, comparing it with several existing RPEs and widely used APEs. Then, we demonstrate its effectiveness in encoder-only models on long-context retrieval tasks. We also report runtime and memory usage, followed by ablation studies to further validate the contributions of our approach.
@@ 5.2 Bidirectional Models
In our second major experiment, we evaluate the effectiveness of our proposed additive RPE method in learning contextual representations. To do this, we replace the original fixed learnable positional encodings in BERT Devlin et al. (2019) with CABLE during pre-training. We use the 10B-sample FineWeb-Edu dataset and train the models using only the masked language modeling (MLM) objective, following Liu (2019), who showed that removing next sentence prediction (NSP) can improve performance. Our BERT models are based on the bert-base-uncased architecture and are trained on four H100 GPUs with a batch size of 32 and a maximum sequence length of 512 for 14k steps (~1 epoch on FineWeb-Edu-10B). We use Adam Kingma and Ba (2014) with a learning rate of 1e-4. As shown in Figure A, CABLE achieves faster convergence in MLM loss compared to other positional encoding baselines.
Since RPE methods do not impose strict limitations on context length, BERT models trained with these encodings show improved long-context performance compared to most existing encoder-only models. However, standardized long-context benchmarks for encoder-only architectures remain limited. Following Warner et al. (2024), we evaluate long-context performance using the English subset of MLDR Chen et al. (2024), a retrieval benchmark consisting of over 200,000 long documents.
@@
5.3 Runtime and Memory Overhead
Note that training time is measured with batching, while inference is measured without batching by recording the time to sequentially generate a set number of tokens.
@@ 6 Conclusion
We introduced CABLE, a novel relative positional encoding method that learns context-aware biases for each token. This is done by adding learned biases into the attention matrix at each layer of decoder-based transformers. Unlike existing methods such as ALiBi, which rely on constant linear biases, CABLE adapts biases based on the tokens’ roles within the sequence. Our experiments demonstrated that CABLE achieves lower perplexity compared to existing methods and significantly enhances length extrapolation capabilities. Results on the edu-fineweb10B and wikitext-103 datasets revealed that CABLE effectively extrapolates on sequences longer than those seen during training, consistently outperforming other approaches. Ablation studies further validated CABLE’s effectiveness, showing that a model trained with CABLE on sequences of length 1024 can outperform a sinusoidal method trained and tested on sequences of length 8192. These advancements were achieved with only a minimal increase in training time and memory usage, while the increase in inference time was only negligible.
@@ A.1 Bert Models Training
A.2 Results for GPT-Small and Tiny
Tables 3 and 4 present the extrapolation results for GPT-Small and GPT-Tiny models evaluated on the FineWeb-Edu-10B dataset. As shown in Table 3, CABLE consistently outperforms other positional encoding methods across all sequence lengths, particularly at extrapolated lengths beyond 1024 tokens. Both GPT-Small and GPT-Tiny models trained with CABLE achieve significantly lower perplexity than those using ALiBi, RoPE, T5-bias, and other baselines. Notably, standard methods such as sinusoidal or learnable encodings degrade sharply at longer lengths, whereas CABLE maintains stable and superior performance. These results further confirm CABLE’s effectiveness in enhancing length extrapolation, even in smaller model regimes.
| GPT-Small | |||||||||

### v3

@@ Abstract
Transformers often struggle to generalize to longer sequences than those seen during training—a limitation known as length extrapolation. Most existing Relative Positional Encoding (RPE) methods attempt to address this by introducing either fixed linear biases or globally learned biases, which lack the capacity to adapt to different input contexts. In this work, we propose an additive RPE, Context-Aware Biases for Length Extrapolation (CABLE), a method that learns token-specific, context-aware biases for each attention head in transformers. By dynamically adjusting positional biases based on the input sequence, CABLE overcomes the rigidity of fixed RPEs. When evaluated on sequences longer than originally trained with, GPT-2 Medium (334M parameters) with CABLE achieves lower perplexity than counterparts using other widely adopted positional encoding methods. Additionally, by applying CABLE to the BERT base model we improved performance in long-context retrieval tasks. Our method significantly enhances the extrapolation performance of existing RPE methods tested on the FineWeb-Edu-10B and WikiText-103 datasets. Our code is available at: https://github.com/AlgonetLabs/Cable.
@@ 2 Related Work
In this section, we review key approaches to positional encoding, including absolute and relative methods, as well as recent work exploring transformer models without any explicit positional encoding.
@@
Fire Li et al. (2023). The kernel function is defined as , where is an MLP with parameters, is a monotonically increasing function and is a learnable scaler. This formulation allows Fire to assign more attention to distant query-key pairs—contrary to methods like ALiBi, and Kerple, which tend to focus on nearby tokens.
@@
Data-dependent Positional Encoding: Recently, several data-dependent RPEs have been proposed. CoPE Golovneva et al. (2024) applies fixed ALiBi-style biases with a learned binary gate. While effective in its domain (mathematics), it lacks the flexibility of continuous, learnable span control. DAPE Zheng et al. (2024a, b) functions as an augmentation to existing additive positional bias methods and is not a standalone mechanism. Furthermore, its architecture imposes a relatively high computational cost, as it uses feedforward or convolutional layers over the full attention matrix. FoX Lin et al. (2025), a near-concurrent work, introduces forgetting mechanisms while minimizing reliance on positional encodings by learning token-wise biases.
Our method differs from these baselines, as CABLE avoids binarization and instead uses dynamic biases (unlike CoPE), functions as a standalone RPE and applies a lightweight MLP over the input (unlike DAPE), and additionally conditions the bias slope on the query token (unlike FoX).
@@ 4.3 Baselines
RoPE Su et al. (2024): A non-learnable non-additive RPE widely adopted in LLMs such as GPT-2 Brown et al. (2020), LLaMA Touvron et al. (2023), PaLM Chowdhery et al. (2023), and Gemma Team et al. (2024a, b).
ALiBi Press et al. (2021): A non-learnable additive RPE used in models like BLOOM Le Scao et al. (2023) and Falcon Almazrouei et al. (2023).
T5-bias Raffel et al. (2020): A learnable additive RPE used in the T5 model.
Kerple Chi et al. (2022a): A learnable additive RPE with logarithmic and power variants; we use the logarithmic variant due to its superior performance.
Fire Li et al. (2023): A learnable additive RPE designed to give more weight to distant query-key pairs than other methods.
@@ 5 Results
We evaluate the effectiveness of our method across multiple settings. First, we examine its extrapolation capability in decoder-only models for next-token prediction, comparing it against several representative RPEs as well as commonly used APEs. Next, we analyze runtime and memory efficiency, followed by a detailed comparison with another context-aware RPE, namely DAPE Zheng et al. (2024b). We then provide a visualization of the positional biases learned by CABLE. Finally, we conduct an ablation study, demonstrating that certain modifications can further enhance CABLE’s performance in specific scenarios.
@@ 5.1 Length Extrapolation
5.2 Runtime and Memory Overhead
5.3 Data-dependent RPEs: CABLE vs. DAPE
As discussed in the related work section, a few context-aware RPEs have already been proposed, with DAPE Zheng et al. (2024a) and its successor DAPEv2 Zheng et al. (2024b) being the most notable. In this section, we compare CABLE against DAPEv2, since the latter represents a direct improvement over the former.
DAPE and DAPEv2 are data-dependent RPEs that act as augmentations to existing additive positional bias methods rather than standalone mechanisms. However, they come with considerable computational cost: DAPE relies on MLP layers, while DAPEv2 applies convolutional layers over the full attention matrix of shape , leading to higher training and backward-pass overhead. In contrast, CABLE applies a lightweight MLP over the input representations of shape , resulting in significantly lower computational overhead.
Here we augment DAPEv2 with CABLE and Kerple for the comparison. Table 2 presents extrapolation results on the FineWeb-Edu-10B dataset, with models trained at sequence length 512 using GPT-2 Small. The results show that while DAPEv2 improves the performance of both CABLE and Kerple, the overall ranking remains unchanged: CABLE consistently outperforms Kerple, both with and without DAPEv2 augmentation.
Kerple suffers from severe degradation at longer sequence lengths (from 4096 onward), whereas CABLE maintains robust performance even without augmentation. Applying DAPEv2 to kerple can alleviate this weakness, but DAPEv2 introduces substantial computational overhead. For instance, GPT-2 Small with CABLE trains at 1.9M tokens/sec, while DAPEv2+CABLE drops to 0.6M tokens/sec (about 3× slower). Also at inference, DAPEv2+CABLE is 1.7× slower than CABLE alone. Similar slowdowns are observed for Kerple with DAPEv2.
Overall, these findings suggest that CABLE offers the best trade-off between accuracy and efficiency, achieving state-of-the-art extrapolation quality while preserving high training and inference throughput.
| Seq. Len. | CABLE | Kerple | DAPEv2+Cable | DAPEv2+Kerple |
| 512 | 21.17 | 21.41 | 20.41 | 20.56 |
| 1024 | 20.72 | 21.12 | 19.91 | 20.07 |
| 2048 | 20.23 | 22.58 | 19.30 | 19.51 |
| 4096 | 19.60 | 28.04 | 18.55 | 18.79 |
| 8192 | 19.87 | 39.38 | 18.63 | 18.92 |
5.4 Analysis of Positional Biases
In this section, we analyze the positional biases added by CABLE and ALiBi to the original attention scores, with a focus on how contextual information influences these biases. Figure 4 illustrates the relative biases for three different heads in the last layer, alongside the ALiBi biases. Unlike ALiBi, which applies a fixed bias uniformly across all layers and tokens, CABLE biases are both head- and context-dependent. Appendix A.3 presents the complete set of visualizations for all attention heads across all examined methods.
The biases presented in Figure 4 were computed using a GPT-2 Small model trained and evaluated with a sequence length of 512 tokens. The contextual dependence of CABLE can be interpreted as a mechanism that allows the model to dynamically calibrate positional information based on the structure of the input sequence. For example, tokens that serve as anchors (e.g., punctuation, repeated entities, or syntactic markers) may receive weaker biases, enabling the model to maintain stronger attention to them even at long distances. In contrast, less informative tokens may be more heavily penalized. Such adaptive behavior could explain why CABLE provides stronger representational flexibility and, in practice, has been observed to improve extrapolation and performance on tasks requiring nuanced handling of context length. Taken together, these findings suggest that while ALiBi provides a simple and efficient positional prior, its uniformity across layers and tokens constrains its ability to model diverse contextual dependencies. By contrast, CABLE introduces a more expressive bias structure that adapts to both token content and attention head, potentially offering a more faithful integration of positional and semantic information.
@@ 5.5 Bidirectional Models
In this major experiment, we evaluate the effectiveness of our proposed additive RPE method in learning contextual representations. To do this, we replace the original fixed learnable positional encodings in BERT Devlin et al. (2019) with CABLE during pre-training. We use the 10B-sample FineWeb-Edu dataset and train the models using only the masked language modeling (MLM) objective, following Liu (2019), who showed that removing next sentence prediction (NSP) can improve performance. Our BERT models are based on the bert-base-uncased architecture and are trained on four H100 GPUs with a batch size of 32 and a maximum sequence length of 512 for 14k steps (~1 epoch on FineWeb-Edu-10B). We use Adam Kingma and Ba (2014) with a learning rate of 1e-4. As shown in Figure A, CABLE achieves faster convergence in MLM loss compared to other positional encoding baselines.
Since RPE methods do not impose strict limitations on context length, BERT models trained with these encodings show improved long-context performance compared to existing encoder-only models. However, standardized long-context benchmarks for encoder-only architectures remain limited. Following Warner et al. (2024), we evaluate long-context performance using the English subset of MLDR Chen et al. (2024), a retrieval benchmark consisting of over 200,000 long documents.
@@ 6 Conclusion
We introduced CABLE, a novel additive relative positional encoding method that learns context-aware biases for each token by injecting them into the attention matrix at every decoder layer. Unlike constant linear biases in ALiBi, CABLE adapts to tokens’ roles within the sequence. Experiments show that CABLE lowers perplexity, significantly improves length extrapolation, and consistently outperforms baselines on edu-fineweb10B and wikitext-103. Moreover, CABLE improves the long-context retrieval performance of encoder-only models. These gains come with only minimal training overhead but faster inference compared to the existing RPEs.
@@ A.1 Bert Models Training
A.2 Results for GPT-2 Small and Tiny
Tables 4 and 5 present the extrapolation results for GPT-2 Small and GPT-2 Tiny models evaluated on the FineWeb-Edu-10B dataset. As shown in Table 4, CABLE consistently outperforms other positional encoding methods across all sequence lengths, particularly at extrapolated lengths beyond 1024 tokens. Both GPT-2 Small and GPT-2 Tiny models trained with CABLE achieve significantly lower perplexity than those using ALiBi, RoPE, T5-bias, and other baselines. Notably, standard methods such as sinusoidal or learnable encodings degrade sharply at longer lengths, whereas CABLE maintains stable and superior performance. These results further confirm CABLE’s effectiveness in enhancing length extrapolation, even in smaller model regimes.
| GPT-2 Small | |||||||||
@@ A.2 Results for GPT-2 Small and Tiny
A.3 Visualizations
Figures 8 to 14 illustrate the additive biases incorporated into the attention scores ( in Equation 1) for the additive RPEs Cable, DAPEv2+Cable, Kerple, DAPEv2+Kerple, ALiBi, Fire, and T5-bias. The visualizations are derived from the final layer of GPT-2 Small model using a randomly selected input example. Since RoPE does not introduce an explicit additive bias but instead modifies the query and key representations, the corresponding attention scores after this transformation are shown separately in Figure 15.